pipe 1 can fill tank in 4 hours time. pipes 1 and 2 together can fill it in 3 hours. pipes 2 and 3 can fill it in 4 hours. how much time would pipe 2 alonr and how much time would pipe 3 alonr require to fill the tank
Answer by TimothyLamb(4379)

a = V/4
a + b = V/3
b + c + V/4
---
a = b + c
a + b + c = a + a
a + b + c = V/4 + V/4
a + b + c = V/2
---
b = V/3 - a
b = V/3 - V/4
b = V4/12 - V3/12
b = V/12
---
b + c + V/4
c = V/4 - b
c = V/4 - V/12
c = V12/48 - V4/48
c = V/6
---
check against earlier a + b + c = V/2:
a + b + c = V/4 + V/12 + V/6
a + b + c = V72/288 + V24/288 + V48/288
a + b + c = V144/288
a + b + c = V/2
---
answer:
pipe 1 alone fills tank in 4 hours
pipe 2 alone fills tank in 12 hours
pipe 3 alone fills tank in 6 hours
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php